# Robust Genomic Prediction and Heritability Estimation using Density Power Divergence

### v1

@@ 2.3.1 The Minimum DPD Estimation for General set-up
$\int f_{\boldsymbol{\theta}}^{(1+\alpha)}-(1+\frac{1}{\alpha})\frac{1}{n}\sum_{i=1}^{n}f_{\boldsymbol{\theta}}^{\alpha}(\textbf{y}_{i}).$
This leads to robust parameter estimates for $\alpha>0$ with the tuning parameter $\alpha$ controlling the trade-off between the (asymptotic) efficiency under pure data and the robustness under contaminated data. It can be seen that the MDPDE at $\alpha=0$ coincides (in a limiting sense) to the most efficient by highly non-robust maximum likelihood estimator (MLE) as the DPD measure at $\alpha=0$ is indeed the Kullback-Leibler divergence. As $\alpha>0$ increases, the corresponding MDPDE achieves greater robustness with a little loss in its pure data (asymptotic) efficiency; see Basu et al. (2011) [3] for more details about the MDPDE along with its applications under standard model set-ups. The theory of the MDPDE has been extended for the independent but non-homogeneous data by Ghosh and Basu (2013) [10] and subsequently found many applications including regression and mixed models. Whenever our observed data $\textbf{y}_{1},\textbf{y}_{2},\ldots,\textbf{y}_{n}$ are independent but non-homogeneous with each $\textbf{y}_{i}\sim g_{i}$ for all $i=1,\ldots,n$ and $g_{i}$ is being modeled by the parametric density $f_{i}(\cdot,\boldsymbol{\theta})$, the MDPDE of the common unknown parameter $\boldsymbol{\theta}$ can be obtained by minimizing the average divergence between the data points and the respective model densities; this becomes equivalent to the minimization of a simpler objective function given by (Ghosh and Basu, 2013 [10])
@@ 2.3.2 The MDPDE based One-stage Approach for Genomic Prediction
For robust genomic prediction in the one-stage approach, we consider the mathematical model described in (1) with $\boldsymbol{\Sigma}_{g}=\sigma_{g}^{2}\mathbb{I}_{p}$ and propose to estimate the associated parameters robustly using the MDPDE instead of the likelihood based approach described in Section 2.2. From (1), we get that $\boldsymbol{y}_{k}\sim N(\mathbf{Z}_{k}\boldsymbol{\gamma},\mathbf{V}_{k})$ independently for each $k=1,\ldots,r$, where
@@
$H_{n}(\boldsymbol{\theta})=\frac{1}{r}\sum_{k=1}^{r}\frac{1}{(2\pi)^{\alpha}|\mathbf{V}_{k}|^{\alpha/2}}\left[\frac{1}{(\alpha+1)}-\left(1+\frac{1}{\alpha}\right)e^{\frac{\alpha}{2}(\mathbf{y}_{k}-\mathbf{Z}_{k}\boldsymbol{\gamma})^{T}\textbf{V}_{k}^{-1}(\mathbf{y}_{k}-\mathbf{Z}_{k}\boldsymbol{\gamma})}\right].$ (10)
In our implementation, we have minimized the above objective function numerically in R to obtain the MDPDE of model parameter in our numerical exploration presented in this paper. One can alternatively compute these MDPDEs by solving the associated estimating equations obtained as $\frac{\partial{H_{n}(\boldsymbol{\theta})}}{\partial\boldsymbol{\gamma}}=\boldsymbol{0}$ and $\frac{\partial{H_{n}(\boldsymbol{\theta})}}{\partial\boldsymbol{\delta}}=0$ with $\boldsymbol{\delta}=(\sigma_{g}^{2},\sigma_{b}^{2},\sigma_{e}^{2})$. We refer to Saraceno et al. (2020) [25] for more details including the asymptotic and robustness properties of the resulting parameter estimates. Once we have obtained the robust MDPDE of the variance parameters as $\widehat{\sigma}_{g}^{2}$ and $\widehat{\sigma}_{e}^{2}$, we can use them to get an estimate of $\boldsymbol{\Lambda}$ as in the case of ordinary ridge regression, by assuming $\boldsymbol{\Lambda}=\lambda\mathbb{I}$ and estimating $\lambda$ by $\widehat{\lambda}=\frac{\widehat{\sigma}_{e}^{2}}{\widehat{\sigma}_{g}^{2}}$. Then the random effects can be robustly estimated from (4) with $\boldsymbol{\Lambda}=\widehat{\lambda}\mathbb{I}$ and subsequently be used for robust estimation of the breeding values and prediction of phenotype effects following (5) and (6), respectively.
@@ 2.3.3 The MDPDE based Two-stage Approach for Genomic Prediction
$\mathbf{y}=\boldsymbol{\eta}\boldsymbol{\mu}+\mathbf{X}_{b}\mathbf{u}_{b}+\boldsymbol{\epsilon}.$ (11)
As described in Piepho et al. (2012) [23], $\boldsymbol{\eta}$ can be taken in such way that $\boldsymbol{\eta}\cdot\boldsymbol{Z}_{1}=\boldsymbol{Z}$ and $\boldsymbol{\eta}\cdot\boldsymbol{X}_{1}=\boldsymbol{X}_{g}$ so that it may be thought of as the design matrix in this simpler model for the genotypic means $\boldsymbol{\mu}=\boldsymbol{Z}_{1}\boldsymbol{\gamma}+\boldsymbol{X}_{1}\boldsymbol{u}_{g}$. In this stage, the marker effects as well as the genotypic mean are considered as fixed parameters and only the block effects are treated as random effects in the model (11). The adjusted mean ($\boldsymbol{\mu}$) can then be estimated easily from the first stage model (11) by classical likelihood based or suitable robust estimators. Following Lourenco et al. (2020) [19], we would use the robust M-estimators with Huber’s weight function, as proposed and studied in Koller (2013) , in this first stage modeling to estimate $\boldsymbol{\mu}$ and associated variance components $\sigma_{b}^{2}$ and $\sigma_{e}^{2}$, which are computed using the function rlmer from the R-package robustlmm (Koller, 2016 [17]) in our numerical experiments. Once we get the estimated adjusted mean as $\widehat{\boldsymbol{\mu}}$, it is then used as response in the second stage model given by
$\widehat{\boldsymbol{\mu}}=\mathbf{Z}_{1}\boldsymbol{\gamma}+\mathbf{X}_{1}\mathbf{u}_{g}+\widetilde{\boldsymbol{\epsilon}}.$ (12)
Assuming normality of these new error terms, we can now estimate $\boldsymbol{\gamma}$, $\sigma_{g}^{2}$ and the new error variance (say $\widetilde{\sigma}_{e}^{2}$) by a suitable method. We propose to use the robust MDPDE for this purpose which can be obtained by solving the associated objective function as described previously. Once we get the robust estimates of all the parameters, we can obtain the robust estimates of the breeding values and prediction of phenotype effects as before following Eqns. (5) and (6), respectively.
@@ 2.4 Measuring Heritability
Mathematically, one can explain heritability to describe the difference of variation in a given trait that can be attributed to genetic variation. To measure how well the differences in genes account for differences in several kind of phenotypic traits, SNP-based heritability can be measured by the ratio of genetic variance $(\sigma_{g}^{2})$ to the total phenotypic variance as well as effect of environment (e.g., block effect); it lies between 00 to $1$ (Wray et al.,  2008 [31]). An estimate of the heritability of a trait is specific for populations as well as environment and changes with circumstances. Once we get the robust MDPDE of the variance parameters as $\widehat{\sigma}_{g}^{2}$, $\widehat{\sigma}_{b}^{2}$ and $\widehat{\sigma}_{e}^{2}$, we can estimate the heritability following Wray et al. (2008 [31]) as
$\widehat{H}_{p}=\frac{\widehat{\sigma}_{g}^{2}}{\widehat{\sigma}_{g}^{2}+\widehat{\sigma}_{b}^{2}+\widehat{\sigma}_{e}^{2}}.$ (13)
@@ 3.1 Experimental Setup: Artificial Datasets
where $y_{ijk}$ is the phenotypic trait value for $i$-th genotype in the $j$-th block within $k$-th replecate, $\phi$ is the overall mean, $g_{i}=\sum_{l=1}^{p}{z_{il}u_{g_{l}}}$ be the $i$-th genotypic value with $z_{il}$ being the SNP-marker values and $u_{g_{l}}$ being the random effects generated from $N(0,\sigma_{g}^{2})$ independently for all $j=1,\ldots,p$, $b_{jk}$ is random block effect of $j$-th block of $k$-th replicate, computed based on random effects associated with $B$ blocks generated independently from $N(0,\sigma_{b}^{2})$ along with the block design matrix, and $e_{ijk}$ is random error generated from $N(0,\sigma_{e}^{2})$ independently for all $i,j,k$. Each artificial dataset is generated with $k=2$ replicates, $B=39$ blocks and $p=11646$ SNP markers, where 4 blocks contain 17 observations each and remaining 35 blocks contain 18 observations each leading to a total of $N=715$ observations. We also take $\phi=0.05$ and $\sigma_{r}^{2}=0$ but the other parameter values are taken differently for two setups as described below.
  • First setup for one-stage approach: Following Lourenco et al. (2020) [19], we take $\sigma_{b}^{2}=6.27$, $\sigma_{e}^{2}=53.8715$ and $\sigma_{g}^{2}=0.005892$. Also the SNP marker values are generated as $z_{il}\sim Bin(1,0.5)$ independently for all $i,l$.
  • Second setup for two-stage approach: We take $\sigma_{b}^{2}=6.3148$, $\sigma_{e}^{2}=20.5$, $\sigma_{s}^{2}=0.5892$, and $z_{il}\sim Bin(2,0.5)-1$ independently for all $i,l$. Note that the minus 1 transform the binomial random variables so that each SNP marker takes value within {-1,0,1}.
@@ 3.2 Competitive Methods
For one-stage approach, we consider the methods described as RMLA and RMLV in Section 2.2 since they are shown to outperform other ridge regression based method by Hofheinz and Frisch (2014) [15]. However, since our primary focus is on robust inference against data contamination, we also compare our MDPDE based one-stage proposal with robust variants of RMLA and RMLV methods, which we would refer to as Rob-RMLA and Rob-RMLV; they are computed as in RMLA and RMLV, respectivley, but using the Huber’s M-estimation approach (Koller, 2013 [16]) to estimate the parameters of the underlying linear mixed models instead of the likelihood based approach. These robust M-estimators are considered as classical benchmark against our proposed MDPDE based approach and are computed using the function rlmer from the R-package robustlmm (Koller, 2016 [17]). For two-stage approach, we consider two existing robust competitors based on the work of Lourenço et al. (2020) [19] which we refer to as Rob1 and Rob2. Both these methods robustify only the first stage estimation by using the M-estimation approach as described in Section 2.3.3; but Rob1 uses the classical likelihood based estimation in the second stage modeling while Rob2 uses the robust M-estimation approach also in the second stage modelling. Recall that our proposal involves using the novel MDPDE based approach in the second stage modeling, while keeping the the first stage the same as Rob1 and Rob2.
@@ 3.3 Performance Metrics
In order to compare our proposed robust procedures with other existing benchmark methodologies of genomic predictions, we illustrate how close the estimated SNP marker effects ($\widehat{\mathbf{u}}_{g}$) are in comparison to their original values ($\mathbf{u}_{g}$). So, for each artificial datasets, we compute the following two measures to illustrate the accuracy of genomic predictions:
  • Pearsonian correlation coefficient $\widehat{\rho}=\frac{Cov(\mathbf{u}_{g},\widehat{\mathbf{u}}_{g})}{\sqrt{Var(\mathbf{u}_{g})Var(\widehat{\mathbf{u}}_{g})}}$. Note that, the higher the value of $\widehat{\rho}$ higher is the predictive accuracy. This measure has been traditionally used in the related literature on genomic prediction (e.g., Lourenco et al., 2020 [19]).
  • Median absolute deviation (MAD): a robust accuracy measure obtained as the median of the absolute difference between the predicted and the original marker effects. A low value of MAD close to zero indicates good performance. We use this measure to get a robust indication of deviations in the predicted trait values.
Note that the values of original $\mathbf{u}_{g}$ are known for each artificial dataset through our simulations. We compute one value of $\widehat{\rho}$ and MAD for each such dataset and report their average over 100 simulation replications in Tables 12. Further, for each artificial dataset we get one final estimated value of the heritability measures $\widehat{H}_{p}$. So, its performances under different existing and proposed approaches are examined by computing its Mean-square deviations (MSD) against its true value (known for each simulation setups) over 100 simulation replications (Table 3).

### v2

@@ 2.3.1 The Minimum DPD Estimation for General set-up
$\int f_{\boldsymbol{\theta}}^{(1+\alpha)}-\left(1+\frac{1}{\alpha}\right)\frac{1}{n}\sum_{i=1}^{n}f_{\boldsymbol{\theta}}^{\alpha}(\textbf{y}_{i}).$
This leads to robust parameter estimates for $\alpha>0$ with the tuning parameter $\alpha$ controlling the trade-off between the (asymptotic) efficiency under pure data and the robustness under contaminated data. It can be seen that the MDPDE at $\alpha=0$ coincides (in a limiting sense) to the most efficient by highly non-robust maximum likelihood estimator (MLE) as the DPD measure at $\alpha=0$ is indeed the Kullback-Leibler divergence. As $\alpha>0$ increases, the corresponding MDPDE achieves greater robustness with a little loss in its pure data (asymptotic) efficiency; see Basu et al. (2011) [3] for more details about the MDPDE along with its applications under standard model set-ups.
The theory of the MDPDE has been extended for the independent but non-homogeneous data by Ghosh and Basu (2013) [14] and subsequently found many applications including regression and mixed models. Whenever our observed data $\textbf{y}_{1},\textbf{y}_{2},\ldots,\textbf{y}_{n}$ are independent but non-homogeneous with each $\textbf{y}_{i}\sim g_{i}$ for all $i=1,\ldots,n$ and $g_{i}$ is being modeled by the parametric density $f_{i}(\cdot,\boldsymbol{\theta})$, the MDPDE of the common unknown parameter $\boldsymbol{\theta}$ can be obtained by minimizing the average divergence between the data points and the respective model densities; this becomes equivalent to the minimization of a simpler objective function given by (Ghosh and Basu, 2013 [14])
@@ 2.3.2 The MDPDE based One-stage Approach for Genomic Prediction
For robust genomic prediction in the one-stage approach, we consider the mathematical model described in (1) with $\boldsymbol{\Sigma}_{g}=\sigma_{g}^{2}\mathbb{I}_{p}$ and propose to estimate the associated parameters robustly using the MDPDE instead of the likelihood based approach described in Section 2.2. In the context of genomic predictions, from the assumed model (1), we get that the $n_{k}$-vector $\boldsymbol{y}_{k}$ containing the target phenotype measurements from the $k$-th replicate has the distribution $\boldsymbol{y}_{k}\sim N(\mathbf{Z}_{k}\boldsymbol{\gamma},\mathbf{V}_{k})$ independently for each $k=1,\ldots,r$, where
@@
$H_{n}(\boldsymbol{\theta})=\frac{1}{r}\sum_{k=1}^{r}\frac{1}{(2\pi)^{\alpha}|\mathbf{V}_{k}|^{\alpha/2}}\left[\frac{1}{(\alpha+1)}-\left(1+\frac{1}{\alpha}\right)e^{\frac{\alpha}{2}(\boldsymbol{y}_{k}-\mathbf{Z}_{k}\boldsymbol{\gamma})^{T}\textbf{V}_{k}^{-1}(\boldsymbol{y}_{k}-\mathbf{Z}_{k}\boldsymbol{\gamma})}\right].$ (10)
In our implementation, we have minimized the above objective function numerically in R to obtain the MDPDE of model parameter in our numerical exploration presented in this paper. One can alternatively compute these MDPDEs by solving the associated estimating equations obtained as $\frac{\partial{H_{n}(\boldsymbol{\theta})}}{\partial\boldsymbol{\gamma}}=\boldsymbol{0}$ and $\frac{\partial{H_{n}(\boldsymbol{\theta})}}{\partial\boldsymbol{\delta}}=0$ with $\boldsymbol{\delta}=(\sigma_{g}^{2},\sigma_{b}^{2},\sigma_{e}^{2})$. We refer to Saraceno et al. (2020) [28] for more details including the asymptotic and robustness properties of the resulting parameter estimates.
Once we have obtained the robust MDPDE of the variance parameters as $\widehat{\sigma}_{g}^{2}$ and $\widehat{\sigma}_{e}^{2}$, we can use them to get an estimate of $\boldsymbol{\Lambda}$ as in the case of ordinary ridge regression, by assuming $\boldsymbol{\Lambda}=\lambda\mathbb{I}$ and estimating $\lambda$ by $\widehat{\lambda}=\frac{\widehat{\sigma}_{e}^{2}}{\widehat{\sigma}_{g}^{2}}$. Then the random effects can be robustly estimated from (4) with $\boldsymbol{\Lambda}=\widehat{\lambda}\mathbb{I}$ and subsequently be used for robust estimation of the breeding values and prediction of phenotype effects following (5) and (6), respectively.
@@ 2.3.3 The MDPDE based Two-stage Approach for Genomic Prediction
$\boldsymbol{y}=\boldsymbol{\eta}\boldsymbol{\mu}+\mathbf{X}_{b}\mathbf{u}_{b}+\boldsymbol{\epsilon}.$ (11)
As described in Piepho et al. (2012) [26], the matrix $\boldsymbol{\eta}$ and the vector $\boldsymbol{\mu}$ can be constructed in such way that $\boldsymbol{\eta}\cdot\boldsymbol{Z}_{1}=\boldsymbol{Z}$ and $\boldsymbol{\eta}\cdot\boldsymbol{X}_{1}=\boldsymbol{X}_{g}$ and then $\boldsymbol{\mu}=\boldsymbol{Z}_{1}\boldsymbol{\gamma}+\boldsymbol{X}_{1}\boldsymbol{u}_{g}$ for some suitably defined matrices $\boldsymbol{Z}_{1}$ and $\boldsymbol{X}_{1}$. Note that, $\boldsymbol{\eta}$ may be thought of as the fixed-effect design matrix with associated parameter vector $\boldsymbol{\mu}$ in this simpler first-stage model (11). Here, both the marker effects ($\mathbf{u}_{g}$) and the genotypic mean ($\boldsymbol{\gamma}$) are considered as fixed parameters within the adjusted mean $\boldsymbol{\mu}$ and only the block effects ($\mathbf{u}_{b}$) are treated as random effects in the model (11). Following Lourenco et al. (2020) [21], we would use the robust M-estimators with Huber’s weight function, as proposed and studied in Koller (2013), in this first stage modeling to estimate $\boldsymbol{\mu}$ and associated variance components $\sigma_{b}^{2}$ and $\sigma_{e}^{2}$. These estimates are computed using the function rlmer from the R-package robustlmm (Koller, 2016 [20]) in our numerical experiments.
Once we get the estimated adjusted mean as $\widehat{\boldsymbol{\mu}}$, it is then used as response in the second stage model given by
$\widehat{\boldsymbol{\mu}}=\mathbf{Z}_{1}\boldsymbol{\gamma}+\mathbf{X}_{1}\mathbf{u}_{g}+\widetilde{\boldsymbol{\epsilon}},$ (12)
where $\widetilde{\boldsymbol{\epsilon}}$ denotes the vector of (new) random errors having variance-covariance matrix same as that of the adjusted mean $\boldsymbol{\mu}$ from the first stage analysis. Assuming normality of these new error terms, we can now estimate $\boldsymbol{\gamma}$, $\sigma_{g}^{2}$ and the new error variance parameter (say $\widetilde{\sigma}_{e}^{2}$) by a suitable method. We propose to use the robust MDPDE for this purpose which can be obtained by solving the associated objective function as described previously. Once we get the robust estimates of all the parameters, we can obtain the robust estimates of the breeding values and prediction of phenotype effects as before following Eqns. (5) and (6), respectively.
@@ 2.4 Measuring Heritability
The extent to which genetic differences contribute to the observed variability in a particular trait within a population is measured by SNP-based heritability. Mathematically, it can be measured by the ratio of genetic variance $(\sigma_{g}^{2})$ to the total phenotypic variance ($\frac{\widehat{\sigma}_{e}^{2}}{r}$), so that it lies between 00 to $1$ An estimate of the heritability of a trait is specific for populations as well as environment and changes with circumstances. Once we get the robust MDPDE of the variance parameters as $\widehat{\sigma}_{g}^{2}$ and $\widehat{\sigma}_{e}^{2}$, we can estimate the heritability following as
$\widehat{H}_{p}=\frac{\widehat{\sigma}_{g}^{2}}{\widehat{\sigma}_{g}^{2}+\frac{\widehat{\sigma}_{e}^{2}}{r}}.$ (13)
@@ 3.1 Experimental Setup: Artificial Datasets
where $y_{ijk}$ is the phenotypic trait value for $i$-th genotype in the $j$-th block within $k$-th replicate, $\phi$ is the overall mean, $g_{i}=\sum_{l=1}^{p}{z_{il}u_{g_{l}}}$ be the $i$-th genotypic value with $z_{il}$ being the SNP-marker values and $u_{g_{l}}$ being the random effects generated from $N(0,\sigma_{g}^{2})$ independently for all $j=1,\ldots,p$, $b_{jk}$ is random block effect of $j$-th block of $k$-th replicate, computed based on random effects associated with $B$ blocks generated independently from $N(0,\sigma_{b}^{2})$ along with the block design matrix, and $e_{ijk}$ is random error generated from $N(0,\sigma_{e}^{2})$ independently for all $i,j,k$.
In order to perform the simulations following a realistic genetic architecture, commonly used in plant breeding scenarios, we have used the R-package AlphaSimR [12]. First, we have simulated the haplotype sequence of the founder population using the MaCS (Markovian Coalescent Simulator) program [6] that is implemented within the AlphaSimR package; the parameter setup in simulating the founder population are taken from [21, 23]. In particular, we have taken $N=715$ individuals where each individuals have 10 chromosomes with 11640 segregating sites per chromosome. Subsequently, we have set our global parameters to model a single trait, utilizing the R-function SimParam, assuming 1000 QTLs per chromosome and SNP markers to be evenly distributed across the 10 chromosomes. The heritability value is set at 0.7 resembling the work done in Muleta et al.  [23]. Then, we have used the function newPop to simulate our initial population based on the founder haplotypes. The SNP-marker matrix (with elements $z_{il}$) is generated using the function pullSnpGeno available within the package AlphaSimR; the SNP marker values are kept within $\{0,1\}$, where 00 is kept as homo-zygotes aa and AA, and $1$ is kept for hetero-zygote Aa. Finally, the function bv from the same package AlphaSimR is used to generate our respective breeding values.
Each artificial dataset is generated following this model set-up with $k=2$ replicates, $B=70$ blocks and $p=11646$ SNP markers, where 5 blocks contain 13 observations each and remaining 65 blocks contain 10 observations each, resulting in the total of $N=715$ observations. For simulating the block effects, we have utilized the design.rcbd and design.alpha functions from the R-package Agricolae [7]. The parameter values are taken as considered in the study by Lourenco et al. [21]; in particular, we take $\phi=0.05$, $\sigma^{2}_{r}=0$, $\sigma^{2}_{b}=6.27$,$\sigma^{2}_{e}=53.8715$ and $\sigma^{2}_{g}=0.005892$.
@@ 3.2 Competitive Methods
For one-stage approach, we consider the methods described as RMLA and RMLV in Section 2.2 since they are shown to outperform other ridge regression based method by Hofheinz and Frisch (2014) [18]. However, since our primary focus is on robust inference against data contamination, we also compare our MDPDE based one-stage proposal with robust variants of RMLA and RMLV methods, which we would refer to as Rob-RMLA and Rob-RMLV; they are computed as in RMLA and RMLV, respectivley, but using the Huber’s M-estimation approach (Koller, 2013 [19]) to estimate the parameters of the underlying linear mixed models instead of the likelihood based approach. These robust M-estimators are considered as classical benchmark against our proposed MDPDE based approach and are computed using the function rlmer from the R-package robustlmm (Koller, 2016 [20]).
For two-stage approach, we consider two existing robust competitors based on the work of Lourenço et al. (2020) [21] which we refer to as Rob1 and Rob2. Both these methods robustify only the first stage estimation by using the M-estimation approach as described in Section 2.3.3; but Rob1 uses the classical likelihood based estimation in the second stage modeling while Rob2 uses the robust M-estimation approach also in the second stage modelling. Recall that our proposal involves using the novel MDPDE based approach in the second stage modeling, while keeping the the first stage the same as Rob1 and Rob2.
@@ 3.3 Performance Metrics
In order to compare our proposed robust procedures with other existing benchmark methodologies of genomic predictions, we illustrate how close the predicted values of the phenotype traits ($\widehat{\boldsymbol{y}}$) are in comparison to their original values ($\boldsymbol{y}$). So, for each artificial datasets, we compute the following two measures to illustrate the accuracy of genomic predictions:
  • Pearson correlation coefficient $\widehat{\rho}=\frac{Cov(\boldsymbol{y},\widehat{\boldsymbol{y}})}{\sqrt{Var(\boldsymbol{y})Var(\widehat{\boldsymbol{y}})}}$. Note that, the higher the value of $\widehat{\rho}$ higher is the predictive accuracy. This measure has been traditionally used in the related literature on genomic prediction (e.g., Lourenco et al., 2020 [21]).
  • Median absolute deviation (MAD): a robust accuracy measure obtained as the median of the absolute difference between the predicted and the original values of the target phenotype traits. A low value of MAD close to zero indicates good performance. We use this measure to get a robust indication of deviations in the predicted trait values.
We compute one value of $\widehat{\rho}$ and MAD for each artificial dataset in our simulations using cross-validation, with a random 70-30 split into training and test data, and report their average over 100 simulation replications in Tables 12. These results represent the out-of-sample prediction accuracy of different methods, obtained based on the test data.
Further, for each artificial dataset we get one final estimated value of the heritability measures $\widehat{H}_{p}$. So, its performances under different existing and proposed approaches are examined by computing its Mean-square deviations (MSD) against its true value (known for each simulation setups) over 100 simulation replications (Table 3).